OpenStudy (anonymous):

I have a test tomorrow, please HELP me! How would I tell if a function is odd, even, or neither even nor odd?

OpenStudy (anonymous):

For an example, consider the following function (of course, it passes the vertical line test, but doesnt look like it on the handdrawn sketch: 

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

How can you tell graphically? And looking at a given function?

OpenStudy (accessdenied):

Well, what is the definition of an odd function and an even function?

OpenStudy (anonymous):

I dont know ...

OpenStudy (accessdenied):

Well, an odd function will satisfy: f(-x) = -f(x). On the graph of f(x), this literally means that the values of the graph on the left side of the y-axis (negative values of x) is going to be the negative value for that f(x) for the positive x.

OpenStudy (accessdenied):

For an even function, the condition is that "f(-x) = f(x)." Graphically, this means that the left side is going to be a mirror image of the right side because the y-value is the same for a negative x as a positive x. A good example of an even function is y = x^2. x=-2 and x=2 have the same y-value, 4.

OpenStudy (stacey):

If a function is even, then the graph to the left of the y axis will look like a mirror image of the right side as in the graph below. (And as access denied just stated)